# EgoGen: An Egocentric Synthetic Data Generator

### 16 Jan 2024 v1

@@ Abstract
Understanding the world in first-person view is fundamental in Augmented Reality (AR). This immersive perspective brings dramatic visual changes and unique challenges compared to third-person views. Synthetic data has empowered third-person-view vision models, but its application to embodied egocentric perception tasks remains largely unexplored. A critical challenge lies in simulating natural human movements and behaviors that effectively steer the embodied cameras to capture a faithful egocentric representation of the 3D world. To address this challenge, we introduce EgoGen, a new synthetic data generator that can produce accurate and rich ground-truth training data for egocentric perception tasks. At the heart of EgoGen is a novel human motion synthesis model that directly leverages egocentric visual inputs of a virtual human to sense the 3D environment. Combined with collision-avoiding motion primitives and a two-stage reinforcement learning approach, our motion synthesis model offers a closed-loop solution where the embodied perception and movement of the virtual human are seamlessly coupled. Compared to previous works, our model eliminates the need for a pre-defined global path, and is directly applicable to dynamic environments. Combined with our easy-to-use and scalable data generation pipeline, we demonstrate EgoGen’s efficacy in three tasks: mapping and localization for head-mounted cameras, egocentric camera tracking, and human mesh recovery from egocentric views. EgoGen will be fully open-sourced, offering a practical solution for creating realistic egocentric training data and aiming to serve as a useful tool for egocentric computer vision research. Refer to our project page.
@@ 2 Related Work
Human-Related Simulators and Synthetic Data. Previous works primarily focus on simulating robots [75, 58, 100, 103, 86] and autonomous cars [18, 23, 8, 81]. While some incorporated animated digital humans, like in [23] and [73], these efforts relied on pre-recorded motion sequences. Rendering images of people to train perception models has been widely studied such as [47, 94, 90, 19, 74, 20]. In particular, [105] offers large-scale synthetic data for egocentric camera wearer pose estimation but relies on mocap data, lacking realistic and spontaneous interactions with the digital world. In contrast, EgoGen closes the gap in egocentric synthetic data generation for head-mounted devices. Please refer to Sec. S1 for detailed comparisons.
@@
Egocentric Human Pose Estimation. Estimating 3D bodies from RGB images is widely studied from third-person views [39, 44, 43, 67, 42, 48, 10, 45], and egocentric views [51, 123, 91, 106, 105, 29, 61, 50, 122, 126], mostly requiring expensive real-world data paired with ground truth annotations. Besides RGB images, depth images offer explicit 3D information, mitigating scale and shape ambiguity, with the potential to enable broader AR/VR applications. However, depth-based methods, especially for the egocentric view, are underexplored. Most existing works [59, 117, 53, 32, 92, 121, 109, 76] predict 3D body skeletons without expressive body meshes, struggling with challenges like severe body truncations and scene occlusions typical in egocentric views. Such limited attention mainly stems from the scarcity of data, as obtaining high-quality human mesh annotations for real-world depth images is labor-intensive.
@@ 3.1 Ego-Sensing Driven Motion Primitives
Motion Primitive Environment. We implement a finite-horizon environment based on the generative motion primitive model from [127]. Specifically, a motion primitive is defined as a 0.5-second motion clip containing $T=20$ frames in the canonical coordinate, and each frame contains a single agent representation. The primitive model $\mathcal{P}$ is based on the C-VAE framework [96], which takes the first $T_{s}=2$ frames as the condition, and models a conditional probability of the next $T-T_{s}$ frames. Compared to [127] trained on the AMASS dataset [52] with many sport motion sequences, we train $\mathcal{P}$ using the SAMP dataset [33], which focuses on daily activities, better suited for HMD use cases. Our action space $\mathcal{A}$ is the pretrained 128D latent space of $\mathcal{P}$, and the action $\mathbf{a}_{t}$ can be randomly sampled from $\mathcal{A}$.
@@ 3.3 Compositing Learned Motion Primitives
         execute one action, produce one CAMP;
@@
Although CAMPs are trained solely with static scenes, their direct application to dynamic settings is achieved by decomposing jointly trained CAMPs into individual motion primitives and re-compositing them. Our model demonstrates effective generalization, provided that the egocentric sensing is updated with the most recent obstacle location at each timestep. Furthermore, our model is directly applicable to tasks involving complex interactions with other virtual humans. To synthesize crowd motion (Alg. 1), each virtual human employs the same policy to navigate and avoid others. To a specific virtual human, others are seen as dynamic obstacles, represented by body bounding boxes for avoidance. Acknowledging the inherent delay in human reactions when avoiding dynamic obstacles [4], agents take a single CAMP sequentially, instead of in parallel, i.e. the first agent generates its first CAMP and waits for others to complete their first CAMP before all agents move on to prepare their second CAMP. To ensure successful collision avoidance, the agent’s egocentric sensing is updated before taking a new action. This composition of CAMPs synthesizes emergent multi-human behaviors without multi-agent RL algorithms (see Sec. 5.1), enhancing the generalization and scalability.
@@ 5.3 Ablation Studies
Directly training RL with egocentric depth images is ineffective due to our high-dimensional action space, emphasizing the value of the compact egocentric sensing representation. Training agents with strict penetration constraints in crowded scenes directly can result in exploring unreasonable action subspaces, given its unbounded Gaussian nature, leading to unrealistic human poses, highlighting the effectiveness of our two-stage RL training scheme. Without the attention reward, the virtual human’s capability to attend to a specific direction decreases. All ablation studies are evaluated in Replica. See visuals in Sup. Vid. and Sec. S4.3.
@@ 7 Future Work and Potentials
Egocentric perception tasks are broad. We will explore many other egocentric vision tasks with EgoGen as this area grows rapidly such as social understanding and forecasting.
EgoGen could also benefit human-robot interaction. For example, our generative human motion model and lifelike human appearances can be integrated into Habitat 3.0 [73] to further close the sim2real gap for robotic agents.
@@
In particular, VirtualHome [71, 72] also provides rendered egocentric views from head-mounted cameras. However, their egocentric videos lack fluctuating patterns due to the absence of natural human motion; instead, they display robotic-like patterns as Habitat 2.0 [100]. We advance closer to synthesizing more realistic data for head-mounted devices. In addition, they lack a generative human motion model, whereas ours can generate more diverse human motion and trajectories. A very recent work Habitat 3.0 [73] introduced virtual humans to robotic simulation. However, their human locomotion is synthesized by cyclically replaying a walking motion clip from MoCap data along a pre-calculated path with rigid rotations to transition to the next walking direction. Both VirtualHome and Habitat 3.0 have a limited number of human agents and fall short in representing diverse human characters, with limitations in body shapes, ethnic variation, and clothing options.
@@
As mentioned in Sec. 3.2 in crowded scenes, we employ a two-stage RL training scheme. In stage I, we pretrain the policy with a penetration weight of $w_{r_{pene}}=1$ to more effectively encourage the virtual human to avoid obstacles and explicitly not perform penetration termination. After convergence, in stage II, we proceed to fine-tune the policy with a strict penetration termination using a reduced penetration weight of $w_{r_{pene}}=0.1$. Penetration detection involves considering the maximum number of body vertices in penetration within a motion primitive. An episode is terminated if:
@@
Our PPO implementation is based on Tianshou [112]. We list the hyperparameters of PPO in Tab. S3. $c_{1},c_{2}$ are defined in Sec. 3.2. “Repeat per Collect” is the training iterations with the same collected rollouts.
@@
Figure S8: EgoGen addresses the issue of sparsity by populating the dataset with synthetic images. In Fig. 7(a), the sparsity of real-world mapping images is apparent, where each red object represents a camera and each colored dot represents a triangulated 3D point. After applying EgoGen, mapping images are more densely distributed, resulting in denser 3D triangulated points, as shown in Fig. 7(b). In Fig. 7(c) and Fig. 7(d), we augment 7(a) with the same amount of synthetic images using [62] and EgoGen respectively. EgoGen generates synthetic data with a similar distribution as human trajectories as illustrated in 7(d). Results are visualized using Colmap [87]. Note that we only visualize a subset of cameras here.

### 11 Apr 2024 v2

@@ Abstract
Understanding the world in first-person view is fundamental in Augmented Reality (AR). This immersive perspective brings dramatic visual changes and unique challenges compared to third-person views. Synthetic data has empowered third-person-view vision models, but its application to embodied egocentric perception tasks remains largely unexplored. A critical challenge lies in simulating natural human movements and behaviors that effectively steer the embodied cameras to capture a faithful egocentric representation of the 3D world. To address this challenge, we introduce EgoGen, a new synthetic data generator that can produce accurate and rich ground-truth training data for egocentric perception tasks. At the heart of EgoGen is a novel human motion synthesis model that directly leverages egocentric visual inputs of a virtual human to sense the 3D environment. Combined with collision-avoiding motion primitives and a two-stage reinforcement learning approach, our motion synthesis model offers a closed-loop solution where the embodied perception and movement of the virtual human are seamlessly coupled. Compared to previous works, our model eliminates the need for a pre-defined global path, and is directly applicable to dynamic environments. Combined with our easy-to-use and scalable data generation pipeline, we demonstrate EgoGen’s efficacy in three tasks: mapping and localization for head-mounted cameras, egocentric camera tracking, and human mesh recovery from egocentric views. EgoGen will be fully open-sourced, offering a practical solution for creating realistic egocentric training data and aiming to serve as a useful tool for egocentric computer vision research.
@@ 2 Related Work
Human-Related Simulators and Synthetic Data. Previous works primarily focus on simulating robots [75, 58, 100, 103, 86] and autonomous cars [18, 23, 8, 81]. While some incorporated animated digital humans, like in [23, 5, 73], these efforts relied on pre-recorded motion sequences. Rendering images of people to train perception models has been widely studied such as [47, 94, 90, 19, 74, 20, 5]. In particular, [105] offers large-scale synthetic data for egocentric camera wearer pose estimation but relies on mocap data, lacking realistic and spontaneous interactions with the digital world. In contrast, EgoGen closes the gap in egocentric synthetic data generation for head-mounted devices. Please refer to Sec. S1 for detailed comparisons.
@@
Egocentric Human Pose Estimation. Estimating 3D bodies from RGB images is widely studied from third-person views [39, 44, 43, 67, 42, 48, 10, 45], and egocentric views [51, 125, 91, 106, 105, 29, 61, 50, 124, 128, 109], mostly requiring expensive real-world data paired with ground truth annotations. Besides RGB images, depth images offer explicit 3D information, mitigating scale and shape ambiguity, with the potential to enable broader AR/VR applications. However, depth-based methods, especially for the egocentric view, are underexplored. Most existing works [59, 118, 53, 32, 92, 123, 110, 76] predict 3D body skeletons without expressive body meshes, struggling with challenges like severe body truncations and scene occlusions typical in egocentric views. Such limited attention mainly stems from the scarcity of data, as obtaining high-quality human mesh annotations for real-world depth images is labor-intensive.
@@ 3.1 Ego-Sensing Driven Motion Primitives
Environment. Inspired by [119], we aim to learn a library of composable CAMPs. We implement a finite-horizon environment based on the generative motion primitive model from [129]. Specifically, a motion primitive is defined as a 0.5-second motion clip containing $T=20$ frames in the canonical coordinate, and each frame contains a single agent representation. The primitive model $\mathcal{P}$ is based on the C-VAE framework [96], which takes the first $T_{s}=2$ frames as the condition, and models a conditional probability of the next $T-T_{s}$ frames. Compared to [129] trained on the AMASS dataset [52] with many sport motion sequences, we train $\mathcal{P}$ using the SAMP dataset [33], which focuses on daily activities, better suited for HMD use cases. Our action space $\mathcal{A}$ is the pretrained 128D latent space of $\mathcal{P}$, and the action $\mathbf{a}_{t}$ can be randomly sampled from $\mathcal{A}$.
@@ 3.3 Compositing Learned Motion Primitives
         execute the action that maximizes the expected return, and produce one CAMP;
@@
Although CAMPs are trained solely with static scenes, their direct application to dynamic settings is achieved by decomposing jointly trained CAMPs into individual motion primitives and re-compositing them. Our model demonstrates effective generalization by selecting the next best motion primitive from the learned CAMP library to maximize the expected return, provided that the egocentric sensing is updated with the most recent obstacle location at each timestep. Furthermore, our model is directly applicable to tasks involving complex interactions with other virtual humans. To synthesize crowd motion (Alg. 1), each virtual human employs the same policy to navigate and avoid others. To a specific virtual human, others are seen as dynamic obstacles, represented by body bounding boxes for avoidance. Acknowledging the inherent delay in human reactions when avoiding dynamic obstacles [4], agents take a single CAMP sequentially, instead of in parallel, i.e. the first agent generates its first CAMP and waits for others to complete their first CAMP before all agents move on to prepare their second CAMP. To ensure successful collision avoidance, the agent’s egocentric sensing is updated before taking a new action. This composition of CAMPs synthesizes emergent multi-human behaviors without multi-agent RL algorithms (see Sec. 5.1), enhancing the generalization and scalability.
@@ 5.3 Ablation Studies
Directly training RL with egocentric depth images is ineffective due to our high-dimensional action space, emphasizing the value of the compact egocentric sensing representation. Training agents with strict penetration constraints in crowded scenes directly can result in exploring unreasonable action subspaces, given its unbounded Gaussian nature, leading to unrealistic human poses, highlighting the effectiveness of our two-stage RL training scheme. Without the attention reward, the virtual human’s capability to attend to a specific direction decreases. All ablation studies are evaluated in Replica. See visuals in Supp. Vid. and Sec. S4.3.
@@ 6 Conclusion and Future Work
We will explore many other egocentric vision tasks with EgoGen as this area grows rapidly such as social understanding and forecasting. EgoGen could benefit human-robot interaction, e.g., our generative human motion model and lifelike human appearances can be integrated into [73] to close the sim2real gap for robotic agents further.
Acknowledgements. This work is supported by the SNSF project grant 200021 204840 and the SDSC PhD fellowship. We sincerely thank Linfei Pan, Yunke Ao, Qianli Ma, and Maxime Raafat for the valuable discussions.
@@
In particular, VirtualHome [71, 72] also provides rendered egocentric views from head-mounted cameras. However, their egocentric videos lack fluctuating patterns due to the absence of natural human motion; instead, they display robotic-like patterns as Habitat 2.0 [100]. We advance closer to synthesizing more realistic data for head-mounted devices. In addition, they lack a generative human motion model, whereas ours can generate more diverse human motion and trajectories. A very recent work Habitat 3.0 [73] introduced virtual humans to robotic simulation. However, their human locomotion is synthesized by cyclically replaying a walking motion clip from MoCap data along a pre-calculated path with rigid rotations to transition to the next walking direction. Both VirtualHome and Habitat 3.0 have a limited number of human agents and fall short in representing diverse human characters, with limitations in body shapes, ethnic variation, and clothing options. UnrealEgo [5] is a large-scale naturalistic dataset for egocentric 3D human motion capture, which employs downward-facing cameras and replies on mocap data. It is used to estimate the human body pose of the camera wearer. In contrast, we employ front-facing cameras and perform generative human motion synthesis for autonomous virtual humans.
@@
As mentioned in Sec. 3.2 in the main paper in crowded scenes, we employ a two-stage RL training scheme. In stage I, we pretrain the policy with a penetration weight of $w_{r_{pene}}=1$ to more effectively encourage the virtual human to avoid obstacles and explicitly not perform penetration termination. After convergence, in stage II, we proceed to fine-tune the policy with a strict penetration termination using a reduced penetration weight of $w_{r_{pene}}=0.1$. Penetration detection involves considering the maximum number of body vertices in penetration within a motion primitive. An episode is terminated if:
@@
Our PPO implementation is based on Tianshou [113]. We list the hyperparameters of PPO in Tab. S3. $c_{1},c_{2}$ are defined in Sec. 3.2 in the main paper. “Repeat per Collect” is the training iterations with the same collected rollouts.
@@
S2.4 Qualitative Failure Cases
Figure S2: Failure cases.
When evaluation significantly differs from training, failure cases may occur, including penetrations and unreached goals. See Fig. S2.
@@
To further enhance realism, one may consider various face shapes and physics simulations to place egocentric cameras. We agree that simulating headset placement and shifting can enhance realism. However, first, achieving faithful simulation would involve modeling deformable human facial skin and muscles, which is an ongoing research problem. Second, even with perfect modeling of the camera extrinsic shifting, real-world devices can still have instrumental errors that may offset fine-grained simulation. As an alternative, we considered that people can wear the headset with slightly different placements, and introduced noise augmentation to simulate this (See Sec. S5.2: Egocentric camera tracking). It might be effective to use aleatoric uncertainty to model headset shifting and instrumental errors together in a Bayesian way.
@@
Figure S9: EgoGen addresses the issue of sparsity by populating the dataset with synthetic images. In Fig. 8(a), the sparsity of real-world mapping images is apparent, where each red object represents a camera and each colored dot represents a triangulated 3D point. After applying EgoGen, mapping images are more densely distributed, resulting in denser 3D triangulated points, as shown in Fig. 8(b). In Fig. 8(c) and Fig. 8(d), we augment 8(a) with the same amount of synthetic images using [62] and EgoGen respectively. EgoGen generates synthetic data with a similar distribution as human trajectories as illustrated in 8(d). Results are visualized using Colmap [87]. Note that we only visualize a subset of cameras here.
@@
Impact of Training Schemes.
Table S4: Training scheme comparison of HMR. “-scratch”: training with limited real-world data only. “-ft”: transfer learning. “-Mixed training”: training with mixed real and synthetic data together.
(All units are mm) G-MPJPE $\downarrow$ MPJPE $\downarrow$ PA-MPJPE $\downarrow$ V2V $\downarrow$ Depth-scratch 117.7 82.2 54.1 100.6 Depth-ft 90.7 65.2 47.3 81.0 Depth-Mixed training 99.8 72.2 51.5 90.7 RGB-scratch - 90.7 59.9 102.1 RGB-ft - 85.3 56.2 97.2 RGB-Mixed training - 85.5 57.3 98.2
We adopted transfer learning to improve generalization by transferring knowledge from pretraining on synthetic data and refining features through real-world data finetuning. Mixed training, as shown in Tab. S4, is less effective than “-ft”, but still better than training with limited real data only.